Triangle ABC has its coordinate at A(a,-1),B(8,-1) and C(5,4).If the base is twice the height, find a
Answer by ikleyn(52781)

.

The problem does not disclose which of the three sides of the triangle is called "the base"; 

so I will assume that the base is the horizontal segment AB.



Then the length of the base is  |8-a|  units, while the height is 4 - (-1) = 5 units.


We find "a" from this absolute value equation


    | 8-a | = 2*5,  


which is equivalent to


    | a-8 | = 10.



It has two solutions  a= 18  or  a= -2.      


ANSWER. Possible values of "a" are  -2  or  18.

Solved, answered, explained and completed.